[INSTRM-2309] b2 mangin and detectorbox temperature looks flipped Created: 11/Jun/24  Updated: 12/Jun/24

Status: Open
Project: Instrument control development
Component/s: None
Affects Version/s: None
Fix Version/s: None

Type: Bug Priority: Normal
Reporter: yuki.moritani Assignee: cloomis
Resolution: Unresolved Votes: 0
Labels: None
Remaining Estimate: Not Specified
Time Spent: Not Specified
Original Estimate: Not Specified

Issue Links:
Duplicate
duplicates INSTRM-2131 swap first two temp channels on b2, s... Open
Story Points: 1

 Description   

Looking xcu_b2__vistemps table on archiver, I think values for mangin, detectorbox, and spiderrod is swapped: i.e., mangin looks spiderrod, spiderrod looks detectorbox, and detecotbox looks mangin.

Unnamed: 0 id detectorbox mangin spiderrod spreader frontring detector1 detector2
 0 0 4593750715 280.402893 146.137405 153.264404 135.693298 282.777710 162.680496 162.988007
 1 1 4593699714 280.388214 146.135696 153.260300 135.694397 282.734711 162.679993 162.986893
 2 2 4594019579 280.499908 146.151398 153.283203 135.698502 282.989014 162.680695 162.987595
 3 3 4593774040 280.410706 146.138397 153.266495 135.693604 282.793915 162.680496 162.987503
 4 4 4593791531 280.415009 146.139008 153.267395 135.693298 282.808411 162.679794 162.987106

 



 Comments   
Comment by arnaud.lefur [ 11/Jun/24 ]
(rubin3_ics) alefur@shell-ics:~$ oneCmd.py xcu_b2 temps status
2024-06-11T12:30:28.868 sent xcu_b2 temps status (from alefur_shell_ics)
2024-06-11T12:30:28.886 xcu_b2 i vistemps=275.4443,143.2888,150.4335,133.0566,277.2478,162.6901,162.9803
2024-06-11T12:30:28.886 xcu_b2 : temps=275.4443,143.2888,150.4335,133.0566,277.2478,400.0000,400.0000,400.0000,400.0000,400.0000,162.6901,162.9803
(rubin3_ics) alefur@shell-ics:~$ oneCmd.py xcu_b1 temps status
2024-06-11T12:30:34.333 sent xcu_b1 temps status (from alefur_shell_ics)
2024-06-11T12:30:35.503 xcu_b1 i vistemps=151.0828,275.6105,143.0807,134.9645,277.6646,162.0839,162.9854
2024-06-11T12:30:35.503 xcu_b1 : temps=151.0828,275.6105,143.0807,134.9645,277.6646,400.0000,400.0000,400.0000,400.0000,400.0000,162.0839,162.9854

It looks like the channel are swapped on the hardware side.
Looking at the current values, I think just mangin and detectorBox are swapped but I could be wrong.

Comment by cloomis [ 12/Jun/24 ]

We have been here before with these b2 sensors, INSTRM-2131. Along with the discussion over whether two or three were swapped. Along with a question about b4.

Never followed up with the (hardware?) tests though. I remembered putting a swap in code, but that was for a pair of b2 amps in the ccdActor (INSTRM-1100).

Comment by cloomis [ 12/Jun/24 ]

I think I'd vote for three being wrong: (high, low, middle) vs. (middle, high, low). 143 and 151 are different enough to be distinguished.

Comment by arnaud.lefur [ 12/Jun/24 ]

Yes, sorry. I agree.

Generated at Sun Apr 06 19:46:58 JST 2025 using Jira 8.3.4#803005-sha1:1f96e09b3c60279a408a2ae47be3c745f571388b.